# Continuous topological phase transition between $\mathbb{Z}_{2}$ topologically ordered phases

### v2

@@ Abstract
Since the Ising* lines meet at the multi-critical point, a natural expectation of the effective theory describing the physics near and at the multi-critical point is the O(2) $\phi^{4}$ field theory, given by the Lagrangian $\mathcal{L}=\left(\partial\pmb{\phi}\right)^{2}/2+m^{2}\pmb{\phi}^{2}/2+g\pmb{\phi}^{4}/(4!)+\cdots$, where $\pmb{\phi}=(\phi_{0},\phi_{1})$ is a two-component order parameter field corresponding to the condensation order parameters $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$ and $\langle\pmb{e}^{2}\rangle$. The multi-critical point corresponds to the Wilson-Fisher fixed point of the O(2) $\phi^{4}$ field theory, which is the 3D XY* conformal field theory [68, 69]. We can predict critical exponents of the condensation order parameters and the are law coefficients from the field theory [44]. In Fig. 4c, the area law coefficients match the correlation length critical exponent $\nu_{+}=1/(3-\Delta_{+})$, where $\Delta_{+}$ is the scaling dimension of the relevant field $\pmb{\phi}^{2}=\phi_{0}^{2}+\phi_{1}^{2}$ deriving to the gapped phases along $h_{w}=h_{z}$. Fig. 4d shows that the three condensation order parameters exhibit two distinct critical exponents $\beta_{\phi}=\Delta_{\phi}\nu_{+}$ and $\beta_{-}=\Delta_{-}\nu_{+}$, where $\Delta_{\phi}$ and $\Delta_{-}$ are the scaling dimensions of the fields $\phi_{0}$ (or $\phi_{1}$) and $\phi_{0}^{2}-\phi_{1}^{2}$, which correspond to $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$ (or $\langle\pmb{e}^{2}\rangle$) and $\langle\pmb{m}^{2}\rangle$, separately. The numerical results imply that the multi-critical point belongs to the XY* universality class.
@@
Continuous topological phase transition between the toric code and double semion phases. Finally, we investigate the path $h_{x}=-1,h_{w}=2h_{c}-h_{z}$ [line (ii) in Fig. 3], which traverses the TC phase, the DS phase, and the multi-critical point, and is orthogonal to $h_{w}=h_{z}$ [line (i)]. The area law coefficients in Fig. 4e indicate that $\pmb{e}$ and $\pmb{e}^{3}$ are confined when $h_{w}>h_{z}$ and $\pmb{e}^{n}\pmb{m}^{k}$ with $n,k\in\{1,3\}$ are confined when $h_{w}<h_{z}$, while $\pmb{m}$ and $\pmb{m}^{3}$ are confined as long as $h_{w}\neq h_{z}$. According to the condensation order parameter in Fig. 4f, $\pmb{e}^{2}\pmb{m}^{2}$ condenses when $h_{w}>h_{z}$, while $\pmb{e}^{2}$ condenses when $h_{w}<h_{z}$. By comparing the numerical results in Figs. 4e and f with Tab. 1, we know along the line (ii), it is the TC phase when $h_{z}>h_{w}$ and the DS phase when $h_{w}>h_{z}$.
@@
As shown in Fig. 4g, we find that the area law coefficients exhibit a critical exponent consistent with the correlation length exponent $\nu_{-}=1/(3-\Delta_{-})$, where $\Delta_{-}$ is the scaling dimension of the relevant field $\phi_{0}^{2}-\phi_{1}^{2}$ deriving the multi-critical point to the TC or DS phases. Fig. 4h shows that the condensation order parameter exhibits the critical exponents $\beta_{\phi}^{\prime}=\Delta_{\phi}\nu_{-}$. The distinct critical exponents from the lines (i) and (ii) in Fig. 3 further confirm that the multi-critical point belongs to the XY* universality class, and the detailed field theory analysis is provided in SM [44].
@@
Interestingly, the continuous TC-DS transition exhibits features that resemble deconfined quantum critical points [41, 42]—a direct continuous transition between two phases that spontaneously break different symmetries and exhibit fractionalized excitations—in the following aspects: First, from Fig. 4f, $\langle\pmb{e}^{2}\rangle=0$ in the DS phase while $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle=0$ in the TC phase, implying the gapped TC and DS phases spontaneously break two different 1-form symmetries, $\prod_{e\in C}Z_{e}^{2}$ and $\prod_{e\in\hat{C}}W_{e}^{2}$, respectively. Second, the anyons $\pmb{m}$ and $\pmb{m}^{3}$ are confined in both the TC and DS phases, but they are deconfined at the continuous TC-DS transition point, indicated by the area law coefficient $\alpha(\pmb{m})$ shown in Fig. 4e; since $\pmb{m}\times\pmb{m}=\pmb{m}^{3}\times\pmb{m}^{3}=\pmb{m}^{2}$ is a well-defined anyon in both TC and DS phases, $\pmb{m}$ and $\pmb{m}^{3}$ are fractionalized anyons with respect to the $\mathbb{Z}_{2}$ topological order. A difference between the continuous TC-DS transition point and deconfined quantum critical points is that two broken 1-form symmetries $\prod_{e\in C}Z_{e}^{2}$ and $\prod_{e\in\hat{C}}W_{e}^{2}$ do not have mixed anomalies.
@@
Discussions and outlook. In this letter, we propose a perturbed $\mathbb{Z}_{4}$ QD model to study the topological phase transitions between the TC and DS phases, from which we find that the direct continuous phase transition exists between these two $\mathbb{Z}_{2}$ topologically ordered phases. We identify that the continuous TC-DS transition belongs to the XY* universality class, where deconfined $\mathbb{Z}_{4}$ anyons emerge. Our findings imply that if we consider a model without any explicit $\mathbb{Z}_{4}$ structure but exhibiting a continuous TC-DS transition, fractionalized anyons can emerge at the transition. We emphasize that the TC-DS transition is not a fine-tuned result since it is a continuous transition line in the plane defined by $h_{x}<0.8$ and $h_{z}=h_{w}$, and we further verified this conclusion through the analysis of O(2) field theory [44]. Since the TC and DS phases are dual to two different $\mathbb{Z}_{2}$ symmetry protected topological (SPT) phases in (2+1)(2+1)[70], the multi-critical point can also be interpreted as a topological critical point between two $\mathbb{Z}_{2}$ SPT phases. Our results shed light on constructing a general theoretical framework for continuous topological phase transitions beyond anyon condensation, i.e. considering emergent anyons to reconcile two incompatible topological orders.
@@
Since the TC–DS transition can be simulated on a digital quantum computer [35], it would be interesting to experimentally realize this transition and measure the corresponding anyonic order parameters. In addition, one may consider adding more perturbations like $h_{x}^{\prime}\sum_{e}X_{e}+\text{h.c.}$, $h_{z}^{\prime}\sum_{e}Z_{e}+\text{h.c.}$, to the perturbed $\mathbb{Z}_{4}$ QD model in Eq. (Continuous topological phase transition between $\mathbb{Z}_{2}$ topologically ordered phases). These terms explicitly break the exact 1-form symmetries $\prod_{e\in C}Z_{e}^{2}$, $\prod_{e\in\hat{C}}X_{e}^{2}$, hence the duality transformation and non-local order parameters do not work anymore. In such cases, the Fredenhagen–Marcu order parameter can be used to detect anyon condensation [71, 72, 57], and an adiabatic approach can be used to investigate confinement [73]. Alternatively, the broken exact 1-form symmetry could be restored via quantum error correction [74].
@@
Acknowledgements. The authors gratefully acknowledge valuable discussions with Lei Wang on numerical computations and support in providing computational resources. W.-T. Xu thanks R.-Z. Huang for helpful comments and acknowledges support from the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.

### v3

@@ Abstract
Since the Ising* lines meet at the multi-critical point, a natural expectation of the effective theory describing the physics near and at the multi-critical point is the O(2) $\phi^{4}$ field theory, given by the Lagrangian $\mathcal{L}=\left(\partial\pmb{\phi}\right)^{2}/2+m^{2}\pmb{\phi}^{2}/2+g\pmb{\phi}^{4}/(4!)+\cdots$, where $\pmb{\phi}=(\phi_{0},\phi_{1})$ is a two-component order parameter field corresponding to the condensation order parameters $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$ and $\langle\pmb{e}^{2}\rangle$. From the lattice model perspective, the duality transformation gauges the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry of the two-coupled Ising model, such that the perturbed $\mathbb{Z}_{4}$ QD model is obtained. We should also gauge the $\mathbb{Z}_{2}$ symmetry $\phi_{i}\leftrightarrow-\phi_{i}$ in the O(2) $\phi^{4}$ field to get the effective field theory describing the multi-critical point of the perturbed $\mathbb{Z}_{4}$ QD model, and the gauging orbifolds the $\phi_{i}$ fields, i.e., $\phi_{i}$ can only appear in pairs and $\phi_{i}$ and $\phi_{i}$ become identical [62]. Therefore, the multi-critical point of the two-coupled Ising (perturbed $\mathbb{Z}_{4}$ quantum double) model corresponds to the Wilson-Fisher fixed point of the ungauged (gauged) O(2) $\phi^{4}$ field theory, namely the 3D XY (XY*) conformal field theory [66, 67], see SM [42] for more discussions on the distinctions between the XY and XY* theories. We can predict critical exponents of the condensation order parameters and the area law coefficients from the field theory [42]. In Fig. 4c, the area law coefficients match the correlation length critical exponent $\nu_{+}=1/(3-\Delta_{+})$, where $\Delta_{+}$ is the scaling dimension of the relevant field $\pmb{\phi}^{2}=\phi_{0}^{2}+\phi_{1}^{2}$ deriving to the gapped phases along $h_{w}=h_{z}$. Fig. 4d shows that the three condensation order parameters exhibit two distinct critical exponents $\beta_{\phi}=\Delta_{\phi}\nu_{+}$ and $\beta_{-}=\Delta_{-}\nu_{+}$, where $\Delta_{\phi}$ and $\Delta_{-}$ are the scaling dimensions of the fields $\phi_{0}$ (or $\phi_{1}$) and $\phi_{0}^{2}-\phi_{1}^{2}$, which correspond to $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$ (or $\langle\pmb{e}^{2}\rangle$) and $\langle\pmb{m}^{2}\rangle$, separately. The numerical results imply that the multi-critical point belongs to the XY* universality class.
@@
Continuous topological phase transition between the toric code and double semion phases. Finally, we investigate the path $h_{x}=-1,h_{w}=2h_{c}-h_{z}$ [line (ii) in Fig. 3], which traverses the TC phase, the DS phase, and the multi-critical point, and is orthogonal to $h_{w}=h_{z}$ [line (i)]. The area law coefficients in Fig. 4e indicate that $\pmb{e}$ and $\pmb{e}^{3}$ are confined when $h_{w}>h_{z}$ and $\pmb{e}^{n}\pmb{m}^{k}$ with $n,k\in\{1,3\}$ are confined when $h_{w}<h_{z}$, while $\pmb{m}$ and $\pmb{m}^{3}$ are confined whenever $h_{w}\neq h_{z}$. According to the condensation order parameters in Fig. 4f, $\pmb{e}^{2}\pmb{m}^{2}$ condenses when $h_{w}>h_{z}$, while $\pmb{e}^{2}$ condenses when $h_{w}<h_{z}$. By comparing the numerical results in Figs. 4e and f with Tab. 1, we conclude that along the line (ii), the system is in the TC phase when $h_{z}>h_{w}$ and in the DS phase when $h_{w}>h_{z}$.
@@
As shown in Fig. 4g, we find that the area law coefficients exhibit a critical exponent consistent with the correlation length exponent $\nu_{-}=1/(3-\Delta_{-})$, where $\Delta_{-}$ is the scaling dimension of the relevant field $\phi_{0}^{2}-\phi_{1}^{2}$ deriving the multi-critical point to the TC or DS phases. Fig. 4h shows that the condensation order parameter exhibits the critical exponent $\beta_{\phi}^{\prime}=\Delta_{\phi}\nu_{-}$. The distinct critical exponents from the lines (i) and (ii) in Fig. 3 further confirm that the multi-critical point belongs to the XY* universality class, and a detailed field theory analysis is provided in SM [42].
@@
Interestingly, the continuous TC-DS transition exhibits features that resemble deconfined quantum critical points [39, 40]—a direct continuous transition between two phases that spontaneously break different symmetries and exhibit fractionalized excitations—in the following aspects: First, from Fig. 4f, $\langle\pmb{e}^{2}\rangle=0$ in the DS phase while $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle=0$ in the TC phase, implying the gapped TC and DS phases spontaneously break two distinct 1-form symmetries, $\prod_{e\in C}Z_{e}^{2}$ and $\prod_{e\in\hat{C}}W_{e}^{2}$, respectively. Second, the anyons $\pmb{m}$ and $\pmb{m}^{3}$ are confined in both the TC and DS phases, but they are deconfined at the continuous TC-DS transition point, indicated by the area law coefficient $\alpha(\pmb{m})$ shown in Fig. 4e; since $\pmb{m}\times\pmb{m}=\pmb{m}^{3}\times\pmb{m}^{3}=\pmb{m}^{2}$ is a well-defined anyon in both TC and DS phases, $\pmb{m}$ and $\pmb{m}^{3}$ are fractionalized anyons with respect to the $\mathbb{Z}_{2}$ topological order. A difference between the continuous TC-DS transition point and deconfined quantum critical points is that two broken 1-form symmetries $\prod_{e\in C}Z_{e}^{2}$ and $\prod_{e\in\hat{C}}W_{e}^{2}$ do not have mixed anomalies.
@@
Discussions and outlook. In this letter, we propose a perturbed $\mathbb{Z}_{4}$ QD model to study the topological phase transitions between the TC and DS phases, from which we find that a direct continuous phase transition exists between these two $\mathbb{Z}_{2}$ topologically ordered phases. We identify that the continuous TC-DS transition belongs to the XY* universality class, where deconfined $\mathbb{Z}_{4}$ anyons emerge. Our findings imply that if we consider a model without any explicit $\mathbb{Z}_{4}$ structure but exhibiting a continuous TC-DS transition, fractionalized anyons can emerge at the transition. We emphasize that the TC-DS transition is not a fine-tuned result, since it is a continuous transition line in the plane defined by $h_{x}<0.8$ and $h_{z}=h_{w}$, and we further verified this conclusion through the analysis of O(2) field theory [42]. Since the TC and DS phases are dual to two different $\mathbb{Z}_{2}$ symmetry protected topological (SPT) phases in (2+1)(2+1)D [68], the multi-critical point can also be interpreted as a topological critical point between two $\mathbb{Z}_{2}$ SPT phases. We note that a topological phase transition with an exact continuous symmetry can realize a deconfined quantum critical point [69]. It is therefore interesting to consider a model that exhibits the continuous TC–DS transition while maintaining an exact continuous symmetry at the critical point. Our results shed light on constructing a general theoretical framework for continuous topological phase transitions beyond anyon condensation, i.e. considering emergent anyons to reconcile two incompatible topological orders.
@@
Since the TC–DS transition can be simulated on a digital quantum computer [33], it would be interesting to experimentally realize this transition and measure the corresponding anyonic order parameters. In addition, one may consider adding more perturbations like $h_{x}^{\prime}\sum_{e}X_{e}+\text{h.c.}$, $h_{z}^{\prime}\sum_{e}Z_{e}+\text{h.c.}$ to the perturbed $\mathbb{Z}_{4}$ QD model in Eq. (Continuous topological phase transition between $\mathbb{Z}_{2}$ topologically ordered phases). These terms explicitly break the exact 1-form symmetries $\prod_{e\in C}Z_{e}^{2}$, $\prod_{e\in\hat{C}}X_{e}^{2}$, hence the duality transformation and non-local order parameters do not work anymore. In such cases, the Fredenhagen–Marcu order parameter can be used to detect anyon condensation [70, 71, 55], and an adiabatic approach can be used to investigate confinement [72]. Alternatively, the broken exact 1-form symmetry could be restored via quantum error correction [73].
@@
Acknowledgments. The authors appreciate valuable discussions with L. Wang on numerical computations and support in providing computational resources. W.-T. Xu thanks R.-Z. Huang for helpful comments and acknowledges support from the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.